Grade 11Algebra

∫(ex/x)dx=?pls tell me someone the answer.i think it cant be integrated.

Arijit
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
(e^x)/x has no basic subsidiary - you can attempt incorporation by parts however you won't go anyplace. so mathematicians have made up an uncommon capacity called the exponential basic, Ei(x), which is characterized as: Ei(x) = essential, from - interminability to x, of (e^t/t) dt. You can't settle this unequivocally (scientifically). 
Nonetheless, you can do it regarding power arrangement! 
Review that 
e^x = ∑ xⁿ/n!,so 
e^x/x = 1/x + ∑ xⁿ/(n+1)!. 
In this way, 
∫e^x/x dx = ln|x| + ∑ xⁿ/(n·n!) + C.